OpenStudy (anonymous):

Solve 2x2 - 8x = -7.

OpenStudy (anonymous):

2x2 - 8x - 7 = 0 solve quadratic ... and you should obtain 2 valuees for x :)

OpenStudy (shaik0124):

2x2-8x+7=0

OpenStudy (anonymous):

plus 7 to the right side and using this formula -b +- (square root of b^2 - 4ac)/ 2ac

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

2x^2 -8x +7 =0 right?

OpenStudy (anonymous):

then a =2 b = -8 and c = 7

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

is that 2x^2 or (2x)^2 ?

OpenStudy (anonymous):

my options are -2+square root 2 -2+2squate root 2 2+square root 2/2 or 2 +square root of 2 over 2

OpenStudy (anonymous):

what are you talking about?